# Approximating Maximum Edge 2-Coloring

### v2

@@ Abstract
Approximation Algorithms Edge 2-Coloring Matchings
@@ 1.2 Our results
We first show that Algorithm 1 is a $1.5$-approximation algorithm if after the normalization, the graph contains a perfect matching (Lemma 19). We can ensure this property if we normalize a subcubic graph [^1] Recall that a graph is subcubic if no vertex has a degree larger than three., even if before applying the normalization it did not have a perfect matching.
Theorem 1.1 ()
@@
While the three modifications are relatively simple, proving that they are approximation-preserving is non-trivial. If none of these modifications can be applied (anymore), we call the graph normalized. Our key insight is that the number of colors in an optimal solution of a normalized graph can be bounded from above, this is stated as Lemma 1 and shown in Section 3.
Lemma 1 ()
@@
We note that without normalization, an optimal solution can have $n$ colors (e.g., if $G$ is an $n$-cycle). In order to prove Lemma 1, we use the notion of character graphs introduced by Feng et al. [12]. A character graph of an edge 2-coloring is a graph that contains exactly one edge from each color class. We first show that for a normalized graph we can ensure the existence of a nice character graph, which is a character graph with several useful properties. These properties allow for a counting argument with respect to the number of components in the character graph, which allows us to prove the bound in Lemma 1.
@@ 2 The Algorithm
Lemma 2
Let $G$ be a graph and let $G^{\prime}$ be the graph obtained from $G$ via a valid modification from Definition 2. Given a polynomial time $\alpha$-approximation algorithm for $G^{\prime}$ we can obtain a polynomial time $\alpha$-approximation algorithm for $G$.
Proof
An $\alpha$-approximation algorithm for $G^{\prime}$ produces a coloring $\chi^{\prime}$ with at least $\mathsf{opt}(G^{\prime})/\alpha$ colors. Following Definitions 1 and 2 we know that $\mathsf{opt}(G)\leq\mathsf{opt}(G^{\prime})$ and that we can find in polynomial time a coloring $\chi$ for the graph $G$ that uses at least as many colors as $\chi^{\prime}$. Thus, we can find in polynomial time a coloring $\chi$ for $G$ that has at least $\mathsf{opt}(G)/\alpha$ colors, that is, we obtain an $\alpha$-approximation algorithm. ∎
@@
Let $m$ be the number of adjacent triangles that comprise the cactus $C$ and let $T_{1},T_{2},\dots,T_{m}$ be these triangles. Let $N$ be the set of needles.
@@
Note that in $T_{i}$ there cannot be two distinct novel colors incident to a needle as the common vertex would see at least three colors. In particular this implies that if $T_{i}$ is incident to three needles (and thus the only triangle in $C$), there is at most one novel color in $C$, a contradiction to having at least $m+1$ colors. Otherwise let $\{u,v,w\}$ be the vertices of $T_{i}$, with $u$ and $v$ incident to a needle and $w$ not incident to a needle. If $T_{i}$ is a pseudo needle, we remove $T_{i}$ and its two needles from $G^{\prime}$ and introduce a new vertex $w^{\prime}$ and a needle $ww^{\prime}$ with $\chi(ww^{\prime})=c$. Then the resulting cactus has $m-1$ triangles and at least novel $m$ colors, not counting the possible novel color of $T_{m}$ – a contradiction to $m$ being minimal. We show that we can always turn $T_{i}$ into a pseudo needle, which then implies that $C$ and $\chi$ cannot exist.
@@
If $T_{i}$ has two novel colors $c^{\prime},c^{\prime\prime}$, both must be incident to $w$ as otherwise $u$ or $v$ would see three colors. Let $w^{\prime},w^{\prime\prime}\notin\{u,v\}$ be the other two vertices adjacent to $w$. Then $\chi(ww^{\prime}),\chi(ww^{\prime\prime})\in\{c^{\prime}c^{\prime\prime}\}$; w.l.o.g. $\chi(ww^{\prime})=c^{\prime\prime}$. We change $\chi$ by setting $\chi(uw)=\chi(vw)=\chi(uv)=c^{\prime}$. We thus obtain a feasible coloring with the same number of colors but only one novel color in $T_{i}$, i.e., $T_{i}$ is a pseudo needle. ∎
@@
Proof
Lemmas 3, 3 and 3 describe how Modifications 1–3 are performed, and it is shown that they can be applied in polynomial time when given a leaf, a degree-2 vertex, or a simple cactus respectively. Furthermore, it is clear that if $G$ contains a leaf or a degree-2 vertex, we can find it efficiently. Simple cacti can also be found in polynomial time as follows.
@@
We now complete the proof by showing that we can normalize a graph by applying a sequence of Modifications 1–3 that ends after polynomially many steps. Observe that by applying Modifications 1 and 3, we will reduce the number of edges in $G$, but may create a polynomial amount of degree-$2$ vertices. However, applying Modification 2 reduces the number of degree-$2$ vertices by one while keeping the number of edges the same. Hence by applying Modification 2 first whenever it is possible, we reach a point when no further modifications can be made after polynomially many steps. ∎
@@
Proof
The statement follows by a simple application of the local ratio method. For the non-trivial components, we color the edges of graph $G^{\prime}$ according to the output of A. Algorithm 1 colors all trivial components optimally. Thus for each component we obtain an $\alpha$-approximation (or better) and the overall approximation ratio cannot be worse than $\alpha$. ∎
Due to Lemmas 4 and 5, from now on we can assume that the problem instance is a normalized connected graph $G$ with more than two vertices. We now have a closer look at leaves.
Lemma 6
@@
In $\bar{\chi}$, vertex $u$ only sees one color, which allows us to recolor the edge $uv$ with color $c_{2}$ in $\hat{\chi}$. The resulting coloring $\hat{\chi}$ uses one more color than $\bar{\chi}$, hence the same number of colors as $\chi$. The pendant edge $uv$ has a unique color, and all pendant edges which had a unique color under $\chi$ still do so under $\hat{\chi}$. Since in a normalized graph leaves are unique, also the last claim of the lemma follows. ∎
@@ 3 An upper bound on the optimal solution
To show Lemma 1, we analyze the character graph of the given instance.
@@ 3.1 Preparing the character graph
Definition 3 (Character Graph)
@@
Based on Lemma 7, we can avoid cycles within a character graph.
Lemma 8
@@
Proof
Let $H$ be a character graph of $G$. By definition, each edge of $H$ has a unique color. We claim that each connected component of a character graph can be converted to a path. Suppose $H$ has a cycle $C$. Since the graph is normalized, each vertex in $C$ has degree at least three. According to lemma 7, the cycle $C$ cannot contain a chord in $G$, since a chord would be incident to two inner vertices of $H$. Thus, we consider a vertex $v\notin C$ which is adjacent to a vertex $u\in C$. Let us denote the two neighbors of vertex $u$ in $C$, by $w_{1}$ and $w_{2}$. By definition, the edges $w_{1}u$ and $w_{2}u$ have different colors $c_{1}\neq c_{2}$ and therefore $\chi(vu)\in\{c_{1},c_{2}\}$. W.l.o.g, we assume $\chi(vu)=c_{1}$. We replace $w_{1}u$ by $vu$ in $H$ and we obtain a new character graph $H^{\prime}$, which does not contain the cycle $C$. We iteratively apply this change for all cycles, until $H^{\prime}$ becomes cycle-free and thus the lemma follows. ∎
@@
Note that the (total) scope of a vertex does not depend on the chosen ordering. The scope of a vertex captures a natural sequence of dependencies between edge colors. In the following lemma, we show how to avoid free vertices in the scope of vertices, a property which is important in the proof of Lemma 1.
Lemma 9
@@
Let $c$ be the color of $C$, $w$ the first free vertex reached and let $e=uu^{\prime}$ be the edge of $H$ of color $c$, according to Definition 4.
Thus $e\in E(H)$ and there is an edge $f=ww^{\prime}$ in the scope graph of $v,\kappa$ with $\chi(f)=c$. We add $f$ to $H$ and remove $e$ from $H$, which leads to a new valid character graph. Observe that $w^{\prime}$ can only be an inner vertex if $w^{\prime}\in\{u,u^{\prime}\}$, by Lemma 7. In addition, $w^{\prime}$ is either an end vertex or another free vertex. We have therefore simply extended a path in $H$ or created a new path. (In particular, $f$ cannot close a cycle since $w$ is a free vertex.)
@@
If $u$ is an inner vertex (before modifying $H$), it also becomes an end vertex and we have reduced the number of free vertices. Otherwise, $u$ is an end vertex and becomes a free vertex. However, in that case $u$ is contained in a color class $C^{\prime}$ created earlier in $\operatorname{scope}(v)$ according to Definition 4.
In both cases we therefore make progress: we either directly reduce the number of free vertices or we “move” the free vertex closer to $v$. Eventually, if the considered color class is that determined by $v$, we use that $v$ is an inner vertex and thus cannot be turned into a free vertex. ∎
@@
We now strengthen Lemma 9.
Lemma 10
@@
Proof
We only have to handle condition (ii). The remaining proof is analogous to Lemma 9. Let $c:=\chi(uu^{\prime})$ and suppose that $w$ is a free vertex incident to an edge $e$ colored $c$. We then add $e$ to $E(H)$ and remove $uu^{\prime}$. The removal creates two free vertices. However, since $U\cap U^{\prime}=\emptyset$, we can recursively move the free vertices to previous color classes. If $|U|=1$ or $|U^{\prime}|=1$, we apply Lemma 9. ∎
@@
Proof
The lemma follows directly from Lemma 8 and Lemma 10, noting that in the proof of Lemma 10 we do not introduce new cycles. Furthermore, given a normalized graph $G$ and the coloring provided by Lemma 6, we may assume that $H$ is a nice character graph where all pendant vertices of $G$ are endpoints of paths in $H$. ∎
3.2 The proof of Lemma 1
With the preparation of Section 3.1, in this section we prove our main lemma. See 1
Let $G$ be a normalized problem instance and $H$ a nice character graph of $(G,\chi)$. Let $F$ be the set of free vertices, $T$ the set of end vertices and $I$ the set of inner vertices of $H$. We define $n:=|V(H)|$, $i:=|I|$, $f:=|F|$ and $t:=|T|$, and clearly $n=i+t+f$. We show that there is a mapping $\iota$ from the set $I$ of inner vertices to $T\cup F$ with the property that for each vertex $v$ from $T\cup F$ there is at most one inner vertex mapped to $v$ if $v\in T$ and at most two inner vertices are mapped to $v$ if $v\in F$. Intuitively, we can see $\iota$ as an injective mapping, where each vertex in $F$ is split into two vertices. The reason is that we can see a free vertex as a path of length zero and we count two end vertices for each path. Furthermore, we maintain that $\iota$ never maps an inner vertex to a pendant vertex of $G$. We first show that the mapping implies Lemma 1.
Lemma 12
@@ 3.2 The proof of Lemma 1
Proof
The mapping $\iota$ implies $2f+t-\ell\geq i=n-t-f$ and thus $3f+2t\geq n+\ell$ since $\ell$ out of $t$ end vertices cannot be used for the assignment. Each free vertex and each path is a component of the character graph $H$. Therefore the number of components is $f+t/2$, which is minimized if $f=0$ and $t=(n+\ell)/2$. Thus there are at least $\frac{n+\ell}{4}$ components in $H$ and the number of colors is at most $n-n/4-\ell/4=3n/4-\ell/4$, completing the proof. ∎
@@
To gain some intuition, we first argue how to grow an initial triangle of the cactus (see also Figure 3). There are no degree-two vertices in $G$, therefore $v$ has a neighbor $u^{\prime}$ in $G$ that is not a neighbor of $v$ in $H$. Due to Lemma 7, $u^{\prime}$ is not an inner vertex. We note that $u^{\prime}$ cannot be a free vertex either: if $u^{\prime}$ was a free vertex and we could not map $v$ to $u^{\prime}$, then $u^{\prime}$ would be saturated and $v$ would be the third inner vertex adjacent to $u^{\prime}$. By Lemma 10, however, that would mean $u^{\prime}$ seeing 3 colors, as each of the three inner vertices would be connected to $u^{\prime}$ with different colors, contradicting the feasibility of $\chi$. Hence $u^{\prime}$ is an end vertex.
@@
By Lemma 7, $\chi(u^{\prime}v)\neq\chi(u^{\prime}u^{\prime\prime})$ and $\chi(u^{\prime}\hat{v})\neq\chi(u^{\prime}u^{\prime\prime})$. Therefore $c:=\chi(u^{\prime}v)=\chi(u^{\prime}\hat{v})$ and, again by Lemma 7, $v$ has an incident edge $e$ and $\hat{v}$ has an incident edge $\hat{e}$ in $H$ with $\chi(e)=\chi(\hat{e})=c$, which implies $e=\hat{e}$. Therefore, $v$ and $\hat{v}$ are neighbors in the same path $P^{\prime}$ of $H$; note that $P^{\prime}$ is not necessarily distinct from $P$.
@@
Proof
If $u$ is an inner vertex, the claim is true by definition of inner vertices and Invariant 5 If $u$ is an end vertex, it is incident to an edge $e\in E(H)$, which is not in $E(K)$ if $u\notin V(K)$. Let $\{u^{\prime}\}:=\iota^{-1}(u)$. Since $e$ is not incident to $u^{\prime}$ and $u^{\prime}$ sees two colors, $|\chi(u)|=2$. By Invariant 5, none of these colors are contained in $\chi(K)$. If $u$ is a free vertex, there are two vertices $u^{\prime}\neq u^{\prime\prime}$ with $\iota^{-1}(u)=\{u^{\prime},u^{\prime\prime}\}$. Due to Lemma 10, $u^{\prime}$ and $u^{\prime\prime}$ are not adjacent in $H$. Then the claim follows by applying the argument of the end-vertex case twice. ∎
@@
Case 1: Suppose $u^{\prime}\notin V(K)$. Then $\tilde{u}u^{\prime}\notin E(H)$ since both colors in $\chi(\tilde{u})$ have their edges from $H$ within the cactus, by Invariant 5. The same invariant and Lemma 7 imply that $u^{\prime}$ is not an inner vertex. Among the two triangles incident to $\tilde{u}$, the one added later to $V(K)$ was only added if $\tilde{u}$ cannot be mapped to $u^{\prime}$. Due to Invariants 2 and 3, previous mappings to $u^{\prime}$ do not originate in $V(K)$.
There are two possible situations that can occur: $u^{\prime}$ is either a saturated free vertex or a saturated end vertex. In both cases, it sees two distinct colors by Lemma 13. Note that it does not matter if $u^{\prime}$ was mapped directly or using Algorithm 2, due to Invariants 4, 5 and 8. In both cases $|\chi(\tilde{u})\cup\chi(u^{\prime})|=4$ and thus there is no valid color for $\tilde{u}u^{\prime}$ by Lemma 7.
Case 2: Suppose $u^{\prime}\in V(K)$. Then $u^{\prime}$ cannot be an inner vertex or a degree-$4$ vertex, by Lemma 7. Since by Invariant 7 $V(K)$ does not contain free vertices, it has to be an end vertex of degree $2$ in $V(K)$. However, by Invariants 4, 5 and 8, $u^{\prime}$ then also sees two colors distinct from $\chi(\tilde{u})$. ∎
@@
Proof
By Lemma 14, the degrees of both vertices in $V(K)$ are two. We first argue that $u^{\prime}$ sees two colors. If $u^{\prime}$ is an internal vertex this is true by definition. Otherwise $u^{\prime}$ is an end vertex and therefore incident to one edge from $H$. By Invariant 8, the edge is not in $E(K)$ and thus the claim follows from Invariants 4 and 5. Symmetrically, the same claim holds for $u^{\prime\prime}$. In particular, $u^{\prime}u^{\prime\prime}\in E(H)$ since otherwise $|\chi(u^{\prime})\cup\chi(u^{\prime\prime})|=4$ and Lemma 7 applies.
The vertex $u^{\prime}$ does not have further neighbors since otherwise by Lemma 13, it would see more than two colors. The claim for $u^{\prime\prime}$ follows symmetrically. ∎

### v3

@@ Abstract
Approximation Algorithms and Edge 2-Coloring and Matchings
@@ 1.2 Our results
We first show that Algorithm 1 is a $1.5$-approximation algorithm if after the normalization, the graph contains a perfect matching (Lemma 4.1). We can ensure this property if we normalize a subcubic graph [^1] Recall that a graph is subcubic if no vertex has a degree larger than three., even if before applying the normalization it did not have a perfect matching.
Theorem 1.1.
@@
While the three modifications are relatively simple, proving that they are approximation-preserving is non-trivial. If none of these modifications can be applied (anymore), we call the graph normalized. Our key insight is that the number of colors in an optimal solution of a normalized graph can be bounded from above, this is stated as Lemma 1.4 and shown in Section 3.
Lemma 1.4.
@@
We note that without normalization, an optimal solution can have $n$ colors (e.g., if $G$ is an $n$-cycle). In order to prove Lemma 1.4, we use the notion of character graphs introduced by Feng et al. [12]. A character graph of an edge 2-coloring is a graph that contains exactly one edge from each color class. We first show that for a normalized graph we can ensure the existence of a nice character graph, which is a character graph with several useful properties. These properties allow for a counting argument with respect to the number of components in the character graph, which allows us to prove the bound in Lemma 1.4.
@@ 2 The Algorithm
Lemma 2.3.
Let $G$ be a graph and let $G^{\prime}$ be the graph obtained from $G$ via a valid modification from Definition 2.2. Given a polynomial time $\alpha$-approximation algorithm for $G^{\prime}$ we can obtain a polynomial time $\alpha$-approximation algorithm for $G$.
Proof 2.4.
An $\alpha$-approximation algorithm for $G^{\prime}$ produces a coloring $\chi^{\prime}$ with at least $\mathsf{opt}(G^{\prime})/\alpha$ colors. Following Definitions 2.1 and 2.2 we know that $\mathsf{opt}(G)\leq\mathsf{opt}(G^{\prime})$ and that we can find in polynomial time a coloring $\chi$ for the graph $G$ that uses at least as many colors as $\chi^{\prime}$. Thus, we can find in polynomial time a coloring $\chi$ for $G$ that has at least $\mathsf{opt}(G)/\alpha$ colors, that is, we obtain an $\alpha$-approximation algorithm.
@@
Let $m$ be the number of adjacent triangles that form the cactus $C$ and let $T_{1},T_{2},\dots,T_{m}$ be these triangles. Let $N$ be the set of needles.
@@
Note that in $T_{i}$ there cannot be two distinct novel colors incident to a needle as the common vertex would see at least three colors. In particular, this implies that if $T_{i}$ is incident to three needles (and thus the only triangle in $C$), there is at most one novel color in $C$, a contradiction to having at least $m+1$ colors. Otherwise let $\{u,v,w\}$ be the vertices of $T_{i}$, with $u$ and $v$ incident to a needle and $w$ not incident to a needle. If $T_{i}$ is a pseudo needle, we remove $T_{i}$ and its two needles from $G^{\prime}$ and introduce a new vertex $w^{\prime}$ and a needle $ww^{\prime}$ with $\chi(ww^{\prime})=c$. Then the resulting cactus has $m-1$ triangles and at least novel $m$ colors, not counting the possible novel color of $T_{m}$ – a contradiction to $m$ being minimal. We show that we can always turn $T_{i}$ into a pseudo needle, which then implies that $C$ and $\chi$ cannot exist.
@@
If $T_{i}$ has two novel colors $c^{\prime},c^{\prime\prime}$, both must be incident to $w$ as otherwise $u$ or $v$ would see three colors. Let $w^{\prime},w^{\prime\prime}\notin\{u,v\}$ be the other two vertices adjacent to $w$. Then $\chi(ww^{\prime}),\chi(ww^{\prime\prime})\in\{c^{\prime}c^{\prime\prime}\}$; w.l.o.g. $\chi(ww^{\prime})=c^{\prime\prime}$. We change $\chi$ by setting $\chi(uw)=\chi(vw)=\chi(uv)=c^{\prime}$. We thus obtain a feasible coloring with the same number of colors but only one novel color in $T_{i}$, i.e., $T_{i}$ is a pseudo needle.
@@
Proof 2.8.
Lemmas 2.5, 2.5 and 2.5 describe how Modifications 1–3 are performed, and it is shown that they can be applied in polynomial time when given a leaf, a degree-2 vertex, or a simple cactus respectively. Furthermore, it is clear that if $G$ contains a leaf or a degree-2 vertex, we can find it efficiently. Simple cacti can also be found in polynomial time as follows.
@@
We now complete the proof by showing that we can normalize a graph by applying a sequence of Modifications 1–3 that ends after polynomially many steps. Observe that by applying Modifications 1 and 3, we will reduce the number of edges in $G$, but may create a polynomial amount of degree-$2$ vertices. However, applying Modification 2 reduces the number of degree-$2$ vertices by one while keeping the number of edges the same. Hence by applying Modification 2 first whenever it is possible, we reach a point when no further modifications can be made after polynomially many steps.
@@
Proof 2.10.
The statement follows by a simple application of the local ratio method. For the non-trivial components, we color the edges of graph $G^{\prime}$ according to the output of A. Algorithm 1 colors all trivial components optimally. Thus for each component we obtain an $\alpha$-approximation (or better) and the overall approximation ratio cannot be worse than $\alpha$.
Due to Lemmas 2.7 and 2.9, from now on we can assume that the problem instance is a normalized connected graph $G$ with more than two vertices. We now have a closer look at leaves.
Lemma 2.11.
@@
In $\bar{\chi}$, vertex $u$ only sees one color, which allows us to recolor the edge $uv$ with color $c_{2}$ in $\hat{\chi}$. The resulting coloring $\hat{\chi}$ uses one more color than $\bar{\chi}$, hence the same number of colors as $\chi$. The pendant edge $uv$ has a unique color, and all pendant edges which had a unique color under $\chi$ still do so under $\hat{\chi}$. Since in a normalized graph leaves are unique, also the last claim of the lemma follows.
@@ 3 An upper bound on the optimal solution
To show Lemma 1.4, we analyze the character graph of the given instance.
@@ 3.1 Preparing the character graph
Definition 3.1 (Character Graph).
@@
Based on Lemma 3.2, we can avoid cycles within a character graph.
Lemma 3.3.
@@
Proof 3.4.
Let $H$ be a character graph of $G$. By definition, each edge of $H$ has a unique color. We claim that each connected component of a character graph can be converted to a path. Suppose $H$ has a cycle $C$. Since the graph is normalized, each vertex in $C$ has degree at least three. According to lemma 3.2, the cycle $C$ cannot contain a chord in $G$, since a chord would be incident to two inner vertices of $H$. Thus, we consider a vertex $v\notin C$ which is adjacent to a vertex $u\in C$. Let us denote the two neighbors of vertex $u$ in $C$, by $w_{1}$ and $w_{2}$. By definition, the edges $w_{1}u$ and $w_{2}u$ have different colors $c_{1}\neq c_{2}$ and therefore $\chi(vu)\in\{c_{1},c_{2}\}$. W.l.o.g, we assume $\chi(vu)=c_{1}$. We replace $w_{1}u$ by $vu$ in $H$ and we obtain a new character graph $H^{\prime}$, which does not contain the cycle $C$. We iteratively apply this change for all cycles, until $H^{\prime}$ becomes cycle-free and thus the lemma follows.
@@
Note that the (total) scope of a vertex does not depend on the chosen ordering. The scope of a vertex captures a natural sequence of dependencies between edge colors. In the following lemma, we show how to avoid free vertices in the scope of vertices, a property which is important in the proof of Lemma 1.4.
Lemma 3.6.
@@
Let $c$ be the color of $C$, $w$ the first free vertex reached and let $e=uu^{\prime}$ be the edge of $H$ of color $c$, according to Definition 3.5.
Thus $e\in E(H)$ and there is an edge $f=ww^{\prime}$ in the scope graph of $v,\kappa$ with $\chi(f)=c$. We add $f$ to $H$ and remove $e$ from $H$, which leads to a new valid character graph. Observe that $w^{\prime}$ can only be an inner vertex if $w^{\prime}\in\{u,u^{\prime}\}$, by Lemma 3.2. In addition, $w^{\prime}$ is either an end vertex or another free vertex. We have therefore simply extended a path in $H$ or created a new path. (In particular, $f$ cannot close a cycle since $w$ is a free vertex.)
@@
If $u$ is an inner vertex (before modifying $H$), it also becomes an end vertex and we have reduced the number of free vertices. Otherwise, $u$ is an end vertex and becomes a free vertex. However, in that case $u$ is contained in a color class $C^{\prime}$ created earlier in $\operatorname{scope}(v)$ according to Definition 3.5.
In both cases we therefore make progress: we either directly reduce the number of free vertices or we “move” the free vertex closer to $v$. Eventually, if the considered color class is that determined by $v$, we use that $v$ is an inner vertex and thus cannot be turned into a free vertex.
@@
We now strengthen Lemma 3.6.
Lemma 3.9.
@@
Proof 3.10.
We only have to handle condition (ii). The remaining proof is analogous to Lemma 3.6. Let $c:=\chi(uu^{\prime})$ and suppose that $w$ is a free vertex incident to an edge $e$ colored $c$. We then add $e$ to $E(H)$ and remove $uu^{\prime}$. The removal creates two free vertices. However, since $U\cap U^{\prime}=\emptyset$, we can recursively move the free vertices to previous color classes. If $|U|=1$ or $|U^{\prime}|=1$, we apply Lemma 3.6.
@@
Proof 3.12.
The lemma follows directly from Lemma 3.3 and Lemma 3.9, noting that in the proof of Lemma 3.9 we do not introduce new cycles. Furthermore, given a normalized graph $G$ and the coloring provided by Lemma 2.11, we may assume that $H$ is a nice character graph where all pendant vertices of $G$ are endpoints of paths in $H$.
3.2 The proof of Lemma 1.4
With the preparation of Section 3.1, in this section we prove our main lemma. See 1.4
Let $G$ be a normalized problem instance and $H$ a nice character graph of $(G,\chi)$. Let $F$ be the set of free vertices, $T$ the set of end vertices and $I$ the set of inner vertices of $H$. We define $n:=|V(H)|$, $i:=|I|$, $f:=|F|$ and $t:=|T|$, and clearly $n=i+t+f$. We show that there is a mapping $\iota$ from the set $I$ of inner vertices to $T\cup F$ with the property that for each vertex $v$ from $T\cup F$ there is at most one inner vertex mapped to $v$ if $v\in T$ and at most two inner vertices are mapped to $v$ if $v\in F$. Intuitively, we can see $\iota$ as an injective mapping, where each vertex in $F$ is split into two vertices. The reason is that we can see a free vertex as a path of length zero and we count two end vertices for each path. Furthermore, we maintain that $\iota$ never maps an inner vertex to a pendant vertex of $G$. We first show that the mapping implies Lemma 1.4.
Lemma 3.13.
@@
Proof 3.14.
The mapping $\iota$ implies $2f+t-\ell\geq i=n-t-f$ and thus $3f+2t\geq n+\ell$ since $\ell$ out of $t$ end vertices cannot be used for the assignment. Each free vertex and each path is a component of the character graph $H$. Therefore the number of components is $f+t/2$, which is minimized if $f=0$ and $t=(n+\ell)/2$. Thus there are at least $\frac{n+\ell}{4}$ components in $H$ and the number of colors is at most $n-n/4-\ell/4=3n/4-\ell/4$, completing the proof.
@@
To gain some intuition, we first argue how to grow an initial triangle of the cactus (see also Figure 3). There are no degree-two vertices in $G$, therefore $v$ has a neighbor $u^{\prime}$ in $G$ that is not a neighbor of $v$ in $H$. Due to Lemma 3.2, $u^{\prime}$ is not an inner vertex. We note that $u^{\prime}$ cannot be a free vertex either: if $u^{\prime}$ was a free vertex and we could not map $v$ to $u^{\prime}$, then $u^{\prime}$ would be saturated and $v$ would be the third inner vertex adjacent to $u^{\prime}$. By Lemma 3.9, however, that would mean $u^{\prime}$ seeing 3 colors, as each of the three inner vertices would be connected to $u^{\prime}$ with different colors, contradicting the feasibility of $\chi$. Hence $u^{\prime}$ is an end vertex.
@@
By Lemma 3.2, $\chi(u^{\prime}v)\neq\chi(u^{\prime}u^{\prime\prime})$ and $\chi(u^{\prime}\hat{v})\neq\chi(u^{\prime}u^{\prime\prime})$. Therefore $c:=\chi(u^{\prime}v)=\chi(u^{\prime}\hat{v})$ and, again by Lemma 3.2, $v$ has an incident edge $e$ and $\hat{v}$ has an incident edge $\hat{e}$ in $H$ with $\chi(e)=\chi(\hat{e})=c$, which implies $e=\hat{e}$. Therefore, $v$ and $\hat{v}$ are neighbors in the same path $P^{\prime}$ of $H$; note that $P^{\prime}$ is not necessarily distinct from $P$.
@@
Proof 3.16.
If $u$ is an inner vertex, the claim is true by definition of inner vertices and Invariant 5 If $u$ is an end vertex, it is incident to an edge $e\in E(H)$, which is not in $E(K)$ if $u\notin V(K)$. Let $\{u^{\prime}\}:=\iota^{-1}(u)$. Since $e$ is not incident to $u^{\prime}$ and $u^{\prime}$ sees two colors, $|\chi(u)|=2$. By Invariant 5, none of these colors are contained in $\chi(K)$. If $u$ is a free vertex, there are two vertices $u^{\prime}\neq u^{\prime\prime}$ with $\iota^{-1}(u)=\{u^{\prime},u^{\prime\prime}\}$. Due to Lemma 3.9, $u^{\prime}$ and $u^{\prime\prime}$ are not adjacent in $H$. Then the claim follows by applying the argument of the end-vertex case twice.
@@
Case 1: Suppose $u^{\prime}\notin V(K)$. Then $\tilde{u}u^{\prime}\notin E(H)$ since both colors in $\chi(\tilde{u})$ have their edges from $H$ within the cactus, by Invariant 5. The same invariant and Lemma 3.2 imply that $u^{\prime}$ is not an inner vertex. Among the two triangles incident to $\tilde{u}$, the one added later to $V(K)$ was only added if $\tilde{u}$ cannot be mapped to $u^{\prime}$. Due to Invariants 2 and 3, previous mappings to $u^{\prime}$ do not originate in $V(K)$.
There are two possible situations that can occur: $u^{\prime}$ is either a saturated free vertex or a saturated end vertex. In both cases, it sees two distinct colors by Lemma 3.15. Note that it does not matter if $u^{\prime}$ was mapped directly or using Algorithm 2, due to Invariants 4, 5 and 8. In both cases $|\chi(\tilde{u})\cup\chi(u^{\prime})|=4$ and thus there is no valid color for $\tilde{u}u^{\prime}$ by Lemma 3.2.
Case 2: Suppose $u^{\prime}\in V(K)$. Then $u^{\prime}$ cannot be an inner vertex or a degree-$4$ vertex, by Lemma 3.2. Since by Invariant 7 $V(K)$ does not contain free vertices, it has to be an end vertex of degree $2$ in $V(K)$. However, by Invariants 4, 5 and 8, $u^{\prime}$ then also sees two colors distinct from $\chi(\tilde{u})$.
@@
Proof 3.20.
By Lemma 3.17, the degrees of both vertices in $V(K)$ are two. We first argue that $u^{\prime}$ sees two colors. If $u^{\prime}$ is an internal vertex this is true by definition. Otherwise $u^{\prime}$ is an end vertex and therefore incident to one edge from $H$. By Invariant 8, the edge is not in $E(K)$ and thus the claim follows from Invariants 4 and 5. Symmetrically, the same claim holds for $u^{\prime\prime}$. In particular, $u^{\prime}u^{\prime\prime}\in E(H)$ since otherwise $|\chi(u^{\prime})\cup\chi(u^{\prime\prime})|=4$ and Lemma 3.2 applies.
The vertex $u^{\prime}$ does not have further neighbors since otherwise by Lemma 3.15, it would see more than two colors. The claim for $u^{\prime\prime}$ follows symmetrically.